# FineMolTex: Towards Fine-grained Molecular Graph-Text Pre-training

### v2

@@ Abstract
Understanding molecular structure and related knowledge is crucial for scientific research. Recent studies integrate molecular graphs with their textual descriptions to enhance molecular representation learning. However, they focus on the whole molecular graph and neglect frequently occurring subgraphs, known as motifs,which are essential for determining molecular properties. Without such fine-grained knowledge, these models struggle to generalize to unseen molecules and tasks that require motif-level insights. To bridge this gap, we propose FineMolTex, a novel Fine-grained Molecular graph-Text pre-training framework to jointly learn coarse-grained molecule-level knowledge and fine-grained motif-level knowledge. Specifically, FineMolTex consists of two pre-training tasks: a contrastive alignment task for coarse-grained matching and a masked multi-modal modeling task for fine-grained matching. In particular, the latter predicts the labels of masked motifs and words, leveraging insights from each other, thereby enabling FineMolTex to understand the fine-grained matching between motifs and words. Finally, we conduct extensive experiments across three downstream tasks, achieving up to 230% improvement in the text-based molecule editing task. Additionally, our case studies reveal that FineMolTex successfully captures fine-grained knowledge, potentially offering valuable insights for drug discovery and catalyst design.
1 Introduction
Comprehending molecular structure and related knowledge is pivotal in scientific investigations spanning diverse fields, including chemistry, drug discovery, and materials science [11]. Recent advancements in artificial intelligence and machine learning have yielded promising outcomes for molecule-based tasks such as retrosynthesis [47] and drug discovery [11]. The majority of these studies [17, 7, 24, 38, 1] concentrate on the molecular structure, such as SMILES strings, molecular graphs, and geometric structures. They learn molecular representations under supervised signals such as toxicity level and drug activity. However, this type of supervised learning requires extensive and costly labeling of pre-defined categories, hindering its adoption for unseen categories and tasks.
Fortunately, compared to task-specific labeled data, textual descriptions of molecules are fairly abundant. These descriptions can be found in chemical database annotations, research papers in chemistry and biology, and drug instruction sheets [26], providing general information on molecular usage, efficacy, chemical properties, and even detailed insights into specific functional groups and chemical moieties [16]. Hence, several studies explore molecular structures along with their corresponding descriptions. MoleculeSTM [26] and MoMu [35] align the whole molecular graphs with their textual descriptions employing a contrastive learning approach as shown in Figure 1(a). MolCA [27] further utilizes a cross-modal projector to map the graph embedding space to the input space of the language model. In this way, these studies reduce the reliance on task-specific labels.
However, these approaches primarily focus on the overall structure of the molecule level, failing to capture fine-grained knowledge of the sub-molecule level, such as functional groups. A natural tool to model sub-molecular structures is the motif [55], which refers to frequently recurring, significant subgraphs within molecular graphs. Motifs often play a key role in determining the properties of the whole molecular graph [55], and motif-level knowledge is frequently depicted in textual descriptions. As shown in Figure 1(b), a benzene ring indicating aromaticity property is reflected by the mention of “aromatic”, and a carboxyl group is reflected by its name “carboxyl” in the description, revealing a fine-grained matching between motifs and texts.
Modeling the fine-grained motif-level knowledge is crucial for two reasons. First, motif-level knowledge is necessary for the generalization to unseen molecules, which are still largely composed of various motifs that have been seen before. For example, consider the zero-shot graph-text retrieval task shown in Figure 1(c), which aims to find the molecule most relevant to the given text. Even if the model has not been trained on the candidate molecules, it has seen many of the motifs within the unseen molecules such as the benzene and the ethyl group, corresponding to the words “benzene” and “carboxylic”, respectively. Thus, the model can easily recognize the relevant molecule. Second, it bridges the gap for downstream tasks that require fine-grained knowledge. For example, in the molecule editing task illustrated in Figure 1(c), the model aims to modify part of the molecular structure based on textual instruction. This requires the model to understand the names or properties of the motifs like “chloride”.
@@ 1 Introduction
Figure 1: Comparison of molecule- and motif-level learning, and illustration of downstream tasks.
@@
Despite the significance of this fine-grained knowledge, it is challenging to jointly learn both molecule- and motif-level knowledge, and also non-trivial to capture fine-grained matching without supervised signals. To overcome these issues, in this work, we propose a novel Fine-grained Molecular graph-Text framework (FineMolTex) to learn fine-grained motif-level knowledge, as well as coarse-grained molecule-level knowledge. First, we extract each motif or word token with an individual embedding for fine-grained knowledge and utilize two global tokens to holistically represent a molecular graph and its corresponding text description as coarse-grained knowledge. Second, to align coarse- and fine-grained matching, we propose two pre-training tasks respectively: the contrastive alignment task based on the global tokens, and the masked multi-modal learning task based on the motif and word tokens. Specifically, in the masked multi-modal modeling task illustrated in Figure 1(b), we randomly mask some motif and word tokens, and further incorporate a cross-attention transformer layer to integrate the embeddings of motifs and words. By predicting the labels of masked motifs and words based on information from each other, the learning of fine-grained alignment knowledge is enhanced. In summary, we outline our contributions as follows:
@@
  • We are the first to reveal that learning fine-grained motif-level knowledge provides key insight for bridging molecular graphs and text descriptions, further empowering the ability to generalize to unseen molecules and tasks.
@@
  • We introduce a novel framework named FineMolTex, consisting of two self-supervised pre-training tasks, to simultaneously learn coarse- and fine-grained knowledge. In particular, the masked multi-modal learning task enhances the prediction for masked tokens leveraging information from the other modality, promoting the learning of fine-grained alignment information.
@@
  • Experimental results across three downstream tasks underscore the effectiveness of FineMolTex, with a notable improvement of up to 230% in the text-based molecule editing task. Furthermore, case studies demonstrate that FineMolTex effectively aligns motifs and words, further facilitating applications such as drug discovery and catalyst design.
2 Related Work
@@
We provide a brief review on molecular multi-modal learning. Prior works predominantly concentrate on modeling the chemical structures such as 1D SMILES [17], 2D molecular graphs [7, 24, 55], and 3D geometric structures [38, 1, 41]. They utilize supervised signals on a predetermined set, and thus cannot generalize to unseen categories without labeled examples. Recently, KV-PLM [53] bridges this gap by linking SMILES with biomedical texts through a unified language modeling framework. Nonetheless, 1D SMILES may omit certain structural details and fail to identify structural similarities among molecules due to its non-uniqueness. To address these limitations, MoleculeSTM [26] and MoMu [35] employ a contrastive learning approach to align the molecular graph with its corresponding text, thus performing well on unseen molecules and texts. However, these models are less effective on molecule-to-text generation tasks because language models are not yet well-versed in interpreting graphs as generative conditions. Therefore, MolCA [27] introduces a cross-modal projector to align the embedding space of the molecular graph with the language model’s input space, enabling the comprehension of 2D graphs as generative conditions. This approach has also been extended to 3D graph structures, where 3D-MoLM [22] uses a cross-modal projector to synchronize the embedding space of the 3D geometric structure with that of the language model. Additionally, various efforts have been devoted to tackling specific molecular tasks based on textual data, including zero-shot instruction molecular learning [56], molecular reaction prediction [34], and molecular relational modeling [9].
@@
3 The Proposed Approach
We propose FineMolTex, a novel fine-grained molecular graph-text framework, learning both molecule- and motif-level knowledge. The model architecture is outlined in Figure 2. This section first introduces the key components in the architecture, and then describes the two pre-training tasks.
3.1 Key Components of FineMolTex
@@
Tokenization. As shown in Figure 2, for fine-grained modeling, we fragment the molecular graphs and texts into motif tokens and word tokens. We employ the BRICS [6] algorithm to decompose molecular graphs into chemically significant motifs, followed by a post-processing procedure [55] to consolidate the motif vocabulary. We break the textual description into word tokens following SciBERT [3]. For coarse-grained modeling, the global tokens of molecule and text, <MOL> and <CLS>, are inserted at the beginning of the motif and word sequences, respectively, resulting in the sequences $m_{0},m_{1},\dots,m_{J}$ and $t_{0},t_{1},\dots,t_{D}$, where $J$ and $D$ are the lengths of the sequences.
Figure 2: Architecture of FineMolTex. The input is a graph-text pair with both a molecular structure and a corresponding description. The components in the same color share the same weights.
Graph Encoder. Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$ represent a molecular graph with $N$ atoms, where $\mathcal{V}=\{v_{1},v_{2},\dots,v_{N}\}$ is the set of atoms, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ denotes the bonds, and $\mathbf{X}=[\mathbf{x_{1},x_{2},\dots,x_{N}}]\in\mathbb{R}^{N\times\zeta}$ is the atom feature matrix. Here, $\mathbf{x_{i}}$ is the feature vector of atom $v_{i}$, and $\zeta$ is the dimension of atom features. We utilize GraphMVP [25], a pre-trained Graph Isomorphism Network (GIN), to encode each motif token. GraphMVP employs multi-view pre-training to connect 2D topologies and 3D geometries, leveraging the GEOM dataset [2], which contains 250K molecular conformations. Denoting the GraphMVP encoder as $f_{\text{GraphMVP}}$, we encode each atom $v$ into an embedding as follows:
$\mathbf{g_{v}}=f_{\text{GraphMVP}}(\mathbf{x_{v}},\mathbf{x_{u}}),\;u\in\mathcal{N}(v),$ (1)
@@
$\mathbf{g_{\mathcal{G}}}=\text{READOUT}(\{\mathbf{g_{v}}|v\in\mathcal{G}\}),\quad\text{for}\quad\mathcal{G}\in\{m_{1},m_{2},\dots,m_{J}\},$ (2)
@@
To preserve the intrinsic connectivity of the motifs in the original molecule, we generate position embeddings based on the breadth-first search order and incorporate them into the motif embeddings $\mathbf{g_{m_{0}},g_{m_{1}},\dots,g_{m_{J}}}$, resulting in updated embeddings $\mathbf{g^{\prime}_{m_{0}},g^{\prime}_{m_{1}},\dots,g^{\prime}_{m_{J}}}$. Text Encoder. We use SciBERT [3], which has been pre-trained on texts from the chemical and biological domains, as our text encoder, denoted as $f_{\text{bert}}$. It can encode a text sequence as:
$\mathbf{b_{t_{0}},b_{t_{1}},\dots,b_{t_{D}}}=f_{\text{bert}}(t_{0},t_{1},\dots,t_{D}).$ (3)
Subsequently, we add the position embeddings to the token embeddings following previous work [3], yielding $\mathbf{b^{\prime}_{t_{0}},b^{\prime}_{t_{1}},\dots,b^{\prime}_{t_{D}}}$.
Transformer Layer. To capture the contextual information for each token, we use “encoder-style” Transformer layers [39], which consist of a multi-head self-attention layer and a fully connected feed-forward network. This enables the tokens to gather information from other tokens in the same modality. We utilize $f_{\text{trm}_{T}}$ and $f_{\text{trm}_{M}}$ for the text and molecule modality, respectively, as follows.
$\mathbf{z_{t_{0}},z_{t_{1}},\dots,z_{t_{D}}}$ $=f_{\text{trm}_{T}}(\mathbf{b^{\prime}_{t_{0}},b^{\prime}_{t_{1}},\dots,b^{\prime}_{t_{D}}}),$ (4)
$\mathbf{z_{m_{0}},z_{m_{1}},\dots,z_{m_{J}}}$ $=f_{\text{trm}_{M}}(\mathbf{g^{\prime}_{m_{0}},g^{\prime}_{m_{1}},\dots,g^{\prime}_{m_{J}}}).$ (5)
@@
$\text{Attention}(Q_{m},K_{t},V_{t})=\text{softmax}\left(\frac{Q_{m}K_{t}^{T}}{\sqrt{d_{k}}}\right)V_{t},$ (6)
@@
$\mathbf{h_{t_{0}},h_{t_{1}},\dots,h_{t_{D}}}$ $=f_{\text{crs}_{T}}(\mathbf{z_{t_{0}},z_{t_{1}},\dots,z_{t_{D}}}),$ (7)
$\mathbf{h_{m_{0}},h_{m_{1}},\dots,h_{m_{J}}}$ $=f_{\text{crs}_{M}}(\mathbf{z_{m_{0}},z_{m_{1}},\dots,z_{m_{J}}}).$ (8)
3.2 Pre-training Tasks

### v5

# Advancing Molecular Graph-Text Pre-training via Fine-grained Alignment

@@ Abstract.
Understanding molecular structure and related knowledge is crucial for scientific research. Recent studies integrate molecular graphs with their textual descriptions to enhance molecular representation learning. However, they focus on the whole molecular graph and neglect frequently occurring subgraphs, known as motifs, which are essential for determining molecular properties. Without such fine-grained knowledge, these models struggle to generalize to unseen molecules and tasks that require motif-level insights. To bridge this gap, we propose FineMolTex, a novel Fine-grained Molecular graph-Text pre-training framework to jointly learn coarse-grained molecule-level knowledge and fine-grained motif-level knowledge. Specifically, FineMolTex consists of two pre-training tasks: a contrastive alignment task for coarse-grained matching and a masked multi-modal modeling task for fine-grained matching. In particular, the latter predicts the labels of masked motifs and words, which are selected based on their importance. By leveraging insights from both modalities, FineMolTex is able to understand the fine-grained matching between motifs and words. Finally, we conduct extensive experiments across three downstream tasks, achieving up to 230% improvement in the text-based molecule editing task. Additionally, our case studies reveal that FineMolTex successfully captures fine-grained knowledge, potentially offering valuable insights for drug discovery and catalyst design.
Graph Neural Networks, Molecular Graph Pre-training
Part of the work was done as a visiting research student at Singapore Management University.
Corresponding authors.
The code is publicly available at  https://github.com/liushiliushi/FineMolTex.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadaisbn: 979-8-4007-1454-2/25/08doi: 10.1145/3711896.3736834
KDD Availability Link:
The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.15501037.
1. Introduction
@@
Figure 1. Comparison of molecule- and motif-level learning, and illustration of downstream tasks.
Comprehending molecular structure and related knowledge is pivotal in scientific investigations spanning diverse fields, including chemistry, drug discovery, and materials science (Gilmer et al., 2017). Recent advancements in artificial intelligence have yielded promising outcomes for molecule-based tasks such as retrosynthesis (Yan et al., 2020) and drug discovery (Gilmer et al., 2017). The majority of these studies (Krenn et al., 2020; Duvenaud et al., 2015; Liu et al., 2019a; Toshev et al., 2023; Atz et al., 2021) concentrate solely on the molecular structure, such as SMILES strings, molecular graphs, and geometric structures. They learn molecular representations under supervised signals such as toxicity level and drug activity. However, this supervised learning requires extensive and costly labeling of pre-defined categories, limiting the application of previous methods to unseen categories and tasks.
Fortunately, compared to task-specific labeled data, textual descriptions of molecules are fairly abundant. These descriptions can be found in chemical database annotations, research papers in chemistry and biology, and drug instruction sheets (Liu et al., 2023c), providing general information on molecular usage, efficacy, chemical properties, and even detailed insights into specific functional groups and chemical moieties (Kim et al., 2021). Hence, several studies explore molecular structures along with their corresponding descriptions. MoleculeSTM (Liu et al., 2023c) and MoMu (Su et al., 2022) align the whole molecular graphs with their textual descriptions employing a contrastive learning approach, as shown in Figure 1(a). MolCA (Liu et al., 2023a) further utilizes a cross-modal projector to map the graph embedding space to the input space of the language model. In this way, these studies reduce the reliance on task-specific labels.
However, these approaches primarily focus on the overall structure of the molecule level, failing to capture fine-grained knowledge of the sub-molecule level, such as functional groups. A natural tool to model sub-molecular structures is the motif (Zhang et al., 2021), which refers to frequently recurring, significant subgraphs within molecular graphs. Motifs often play a key role in determining the properties of the whole molecular graph (Zhang et al., 2021), and motif-level knowledge is frequently depicted in textual descriptions. As shown in Figure 1(b), a benzene ring indicating aromaticity property is reflected by the mention of “aromatic”, and a carboxyl group is reflected by its name “carboxyl” in the description, revealing a fine-grained matching between motifs and texts.
Modeling the fine-grained motif-level knowledge is crucial for two reasons. First, motif-level knowledge is necessary for the generalization to unseen molecules, which are still largely composed of various motifs that have been seen before. For example, consider the zero-shot graph-text retrieval task shown in Figure  1(c), which aims to find the molecule most relevant to the given text. Even if the model has not been trained on the candidate molecules, it has seen many of the motifs within the unseen molecules such as the benzene and the ethyl group, corresponding to the words “benzene” and “carboxylic”, respectively. Thus, the model can easily recognize the relevant molecule. Second, it bridges the gap for downstream tasks that require fine-grained knowledge. For example, in the molecule editing task illustrated in Figure  1(c), the model aims to modify part of the molecular structure based on textual instruction. This requires the model to understand the names or properties of the motifs like “chloride”.
@@
Despite the significance of this fine-grained knowledge, it is challenging to jointly learn both molecule- and motif-level knowledge, and also non-trivial to capture fine-grained matching without supervised signals. To overcome these issues, in this work, we propose a novel Fine-grained Molecular graph-Text framework (FineMolTex) to learn fine-grained motif-level knowledge, as well as coarse-grained molecule-level knowledge. First, to jointly capture both molecule- and motif-level knowledge, we use motif and word tokens to capture fine-grained knowledge, as well as two global tokens, one for the molecular graph and one for its corresponding text, to capture coarse-grained knowledge. To align this knowledge, we introduce two pre-training tasks: contrastive alignment based on global tokens and masked multi-modal learning based on motif or word tokens. Second, to capture fine-grained matching without supervised signals, as illustrated in Figure 1(b), we propose importance scores to identify important motifs and word tokens that contain crucial fine-grained knowledge, and then selectively mask these tokens based on their importance. Finally, we incorporate a cross-attention transformer layer to integrate the embeddings of motifs and words. By predicting the labels of masked motifs and words based on information from each other, the learning of fine-grained alignment knowledge is enhanced. In summary, we outline our contributions as follows.
@@
  • We reveal that learning fine-grained motif-level knowledge provides key insight for bridging molecular graphs and text descriptions, further empowering the ability to generalize to unseen molecules and tasks.
@@
  • We introduce a novel framework named FineMolTex, consisting of two self-supervised pre-training tasks, to simultaneously learn coarse- and fine-grained knowledge. In particular, the masked multimodal learning task enhances the prediction for important masked tokens leveraging information from the other modality, promoting the learning of fine-grained alignment information.
@@
  • Experimental results across three downstream tasks underscore the effectiveness of FineMolTex, with a notable improvement of up to 238% in the text-based molecule editing task. Furthermore, case studies demonstrate that FineMolTex effectively aligns motifs and words, further facilitating applications such as drug discovery and catalyst design.
2. Related Work
@@
We provide a brief review on molecular multi-modal learning. Prior works predominantly concentrate on modeling the chemical structures such as 1D SMILES (Krenn et al., 2020), 2D molecular graphs (Duvenaud et al., 2015; Liu et al., 2019a; Zhang et al., 2021), and 3D geometric structures (Toshev et al., 2023; Atz et al., 2021; Wang et al., 2023). They utilize supervised signals on a predetermined set, and thus cannot generalize to unseen categories without labeled examples. Recently, KV-PLM (Zeng et al., 2022) bridges this gap by linking SMILES with biomedical texts through a unified language modeling framework. Nonetheless, 1D SMILES may omit certain structural details and fail to identify structural similarities among molecules due to its non-uniqueness. To address these limitations, MoleculeSTM (Liu et al., 2023c) and MoMu (Su et al., 2022) employ a contrastive learning approach to align the molecular graph with its corresponding text, thus performing well on unseen molecules and texts. However, these models are less effective on molecule-to-text generation tasks because language models are not yet well-versed in interpreting graphs as generative conditions. Therefore, MolCA (Liu et al., 2023a) introduces a cross-modal projector to align the embedding space of the molecular graph with the language model’s input space, enabling the comprehension of 2D graphs as generative conditions. This approach has also been extended to 3D graph structures, where 3D-MoLM (Li et al., 2024) uses a cross-modal projector to synchronize the embedding space of the 3D geometric structure with that of the language model. Beyond molecular graphs and textual descriptions, MV-Mol (Luo et al., 2024) further incorporates knowledge graphs, expanding the multi-modal framework to include additional sources. Additionally, various efforts have been devoted to tackling specific molecular tasks based on textual data, including zero-shot instruction molecular learning (Zhao et al., 2023), molecular reaction prediction (Shi et al., 2023), and molecular relational modeling (Fang et al., 2024).
@@
3. The Proposed Approach
Figure 2. Architecture of FineMolTex. The input is a graph-text pair with both a molecular structure and a corresponding description. The components in the same color share the same weights.
We propose FineMolTex, a novel fine-grained molecular graph-text framework, learning both molecule- and motif-level knowledge. The model architecture is outlined in Figure 2. This section first introduces the key components in the architecture and then describes the two pre-training tasks.
3.1. Key Components of FineMolTex
@@
Tokenization. As shown in Figure 2, for fine-grained modeling, we fragment the molecular graphs and texts into motif tokens and word tokens. We employ the BRICS (Degen et al., 2008) algorithm to transform the molecular graph into a motif tree, and then generate a motif sequence following a breadth-first search order. Then we utilize the post-processing procedure  (Zhang et al., 2021) to consolidate the motif vocabulary. We break the textual description into word tokens using the word tokenizer of SciBERT (Beltagy et al., 2019). For coarse-grained modeling, the global tokens of molecule and text, ¡MOL¿ and ¡CLS¿, are inserted at the beginning of the motif and word sequences, respectively, resulting in the sequences $m_{0},m_{1},\dots,m_{J}$ and $t_{0},t_{1},\dots,t_{D}$, where $J$ and $D$ are the lengths of the sequences.
Graph Encoder. Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$ represent a molecular graph with $N$ atoms, where $\mathcal{V}=\{v_{1},v_{2},\dots,v_{N}\}$ is the set of atoms, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ denotes the bonds, and $\mathbf{X}=[\mathbf{x_{1},x_{2},\dots,x_{N}}]\in\mathbb{R}^{N\times\zeta}$ is the atom feature matrix. Here, $\mathbf{x_{i}}$ is the feature vector of atom $v_{i}$, and $\zeta$ is the dimension of atom features. We utilize GraphMVP (Liu et al., 2019b), a pre-trained Graph Isomorphism Network (GIN), to encode each motif token. GraphMVP employs multi-view pre-training to connect 2D topologies and 3D geometries, leveraging the GEOM dataset (Axelrod and Gómez-Bombarelli, 2022), which contains 250K molecular conformations. Denoting the GraphMVP encoder as $f_{\text{GraphMVP}}$, we encode each atom $v$ into an embedding as follows:
(1) $\mathbf{g_{v}}=f_{\text{GraphMVP}}(\mathbf{x_{v}},\mathbf{x_{u}}),\;u\in\mathcal{N}(v),$
@@
(2) $\mathbf{g_{\mathcal{G}}}=\text{READOUT}(\{\mathbf{g_{v}}|v\in\mathcal{G}\}),\quad\forall\quad\mathcal{G}\in\{m_{1},m_{2},\dots,m_{J}\},$
@@
To preserve the intrinsic connectivity of the motifs in the original molecule, we generate position embeddings based on the breadth-first search order and incorporate them into the motif embeddings $\mathbf{g_{m_{0}},g_{m_{1}},\dots,g_{m_{J}}}$, resulting in updated embeddings $\mathbf{g^{\prime}_{m_{0}},g^{\prime}_{m_{1}},\dots,g^{\prime}_{m_{J}}}$.
Text Encoder. We use SciBERT (Beltagy et al., 2019), which has been pre-trained on texts from the chemical and biological domains, as our text encoder, denoted as $f_{\text{bert}}$. It can encode a text sequence as:
(3) $\mathbf{b_{t_{0}},b_{t_{1}},\dots,b_{t_{D}}}=f_{\text{bert}}(t_{0},t_{1},\dots,t_{D}).$
Subsequently, we add the position embeddings to the token embeddings following previous work (Beltagy et al., 2019), yielding $\mathbf{b^{\prime}_{t_{0}},b^{\prime}_{t_{1}},\dots,b^{\prime}_{t_{D}}}$.
Transformer Layer. To capture the contextual information for each token, we use “encoder-style” Transformer layers (Vaswani et al., 2017), which consist of a multi-head self-attention layer and a fully connected feed-forward network. This enables the tokens to gather information from other tokens in the same modality. We utilize $f_{\text{trm}_{T}}$ and $f_{\text{trm}_{M}}$ for the text and molecule modality, respectively, as follows.
(4) $\mathbf{z_{t_{0}},z_{t_{1}},\dots,z_{t_{D}}}$ $=f_{\text{trm}_{T}}(\mathbf{b^{\prime}_{t_{0}},b^{\prime}_{t_{1}},\dots,b^{\prime}_{t_{D}}}),$
(5) $\mathbf{z_{m_{0}},z_{m_{1}},\dots,z_{m_{J}}}$ $=f_{\text{trm}_{M}}(\mathbf{g^{\prime}_{m_{0}},g^{\prime}_{m_{1}},\dots,g^{\prime}_{m_{J}}}).$
@@
(6) $\text{Attention}(Q_{m},K_{t},V_{t})=\text{softmax}\left(\frac{Q_{m}K_{t}^{T}}{\sqrt{d_{k}}}\right)V_{t},$
@@
(7) $\mathbf{h_{t_{0}},h_{t_{1}},\dots,h_{t_{D}}}$ $=f_{\text{crs}_{T}}(\mathbf{z_{t_{0}},z_{t_{1}},\dots,z_{t_{D}}}),$
(8) $\mathbf{h_{m_{0}},h_{m_{1}},\dots,h_{m_{J}}}$ $=f_{\text{crs}_{M}}(\mathbf{z_{m_{0}},z_{m_{1}},\dots,z_{m_{J}}}).$
3.2. Pre-training Tasks